# Locally Convex Global Loss Network for Decision-Focused Learning

### v2

@@ Appendix C Additional Results
C.1 Fake Target Variants Results for Budget Allocation
Here, we show the results of budget allocation with varying numbers of fake targets. We tested problems having $\{0,5,50,500\}$ fake targets, where the problem becomes more challenging as the fake targets increases. We use 16 samples for surrogate loss models. The results are presents in Table 2. The normalized test regret is lower the better, and we bold-lettered the results for each fake target experiments.
@@ C.2 Sample Size Variants Results
We present our experimental results using different sample sizes of $\{2,4,8,16,32\}$ for each problem and methodology. The results of normalized test regret for inventory stock, budget allocation, and portfolio optimization problem are in Table 3, 4, and 5 respectively. Note that the result of PFL and DFL remain unchanged regardless of the number of samples since they do not sample predictions. We bold-lettered the best performing results for each sample sizes.

### v4

@@ Appendix C Additional Results
C.1 Fake Target Variants Results for Budget Allocation Problem
Here, we show the results of budget allocation with varying numbers of fake targets. We tested problems having $\{0,5,50,500\}$ fake targets, where the problem becomes more challenging as the fake targets increase. We use 16 samples for surrogate loss models. The results are presented in Table 2. The normalized test regret is lower the better, and we bold-lettered the results for each fake target experiment.
@@ C.2 Sample Size Variants Results
We present our experimental results using different sample sizes of $\{2,4,8,16,32\}$ for each problem and methodology. The results of normalized test regret for inventory stock, budget allocation, and portfolio optimization problem are in Table 3, 4, and 5 respectively. Note that the results of PFL and DFL remain unchanged regardless of the number of samples since they do not sample predictions. We bold-lettered the best-performing results for each sample size.